5.8 g of an inorganic metal oxide ore mainly contains 90% of Nickel (III) oxide which decomposes to elemental nickel and oxygen gas. Produced 880 mL of oxygen were collected over water at a temperature of 25ยฐC with a total pressure of 747 mmHg. *****What is the percent yield of decomposition reaction?***** (PH20:23.7 mmHg at 25ยฐC)

1
Expert's answer
2022-04-12T10:23:03-0400

m1(Ni2O3)=m(inorganic metal oxide)*90%

m1(Ni2O3)=5.8g * 0.90 = 5.22g


2Ni2O3 = 4Ni + 3O2

m2(Ni2O3) = (V(O2)*2Mr(Ni2O3)) / 3Vm(O2)

m2(Ni2O3) =( 0.880 * (2*166)) / 3*22.4 = 4.35g


percent yield = (m2/m1)*100% = (4.35/5.22)*100% = 83.33%